OpenStudy (wavykelp):

I need help understanding this in detail!!!!!! Find the rational number between these two numbers: 3/10 and 0.4555555555555555555555 you get the point.

OpenStudy (anonymous):

3/10 = .3 the rational expression of 0.45555555555555555555555555555555 can be found by letting \(x=0.4\bar{5}\) and then \(10x=4.\bar{5}\) so that \(9x=4.1\Rightarrow x = \frac{41}{90}\)

OpenStudy (wavykelp):

ok so if 0.45555= 41/90 thennnn 3/10= 27/90?

OpenStudy (anonymous):

if you want the rational number exactly half way inbetween, simply add the numbers together and divide by 2. yes, 3/10 = 27/90

OpenStudy (wavykelp):

so if 27/90+41/90=68/90 then 68/90 divided by 2 then uhmm...?

OpenStudy (anonymous):

remember that dividing by 2 is the same as multiplying by the reciprocal, namely 1/2.

OpenStudy (anonymous):

in fact, dividing by any number is the same as multiplying by its reciprocal

OpenStudy (wavykelp):

ok so 68/90 divided by 1/2 = 0.377777

OpenStudy (wavykelp):

how exactly do i convert that?

OpenStudy (anonymous):

no, like this... \[\frac{68}{90}\cdot \frac{1}{2}=\frac{68\cdot 1}{90\cdot 2}=\frac{\cancel2\cdot \cancel2\cdot 17}{\cancel2\cdot \cancel2\cdot 45}=\frac{17}{45}\]

OpenStudy (wavykelp):

ok let me do this on paper real quick

OpenStudy (wavykelp):

where do the crossed out two's come from?

OpenStudy (anonymous):

\(68 = 4\cdot 17 = 2\cdot 2\cdot 17\) \(2\cdot 90 = 2\cdot 2\cdot 45\)

OpenStudy (anonymous):

does that make sense?

OpenStudy (wavykelp):

opps sry didnt men to do that

OpenStudy (wavykelp):

ya that makes sense thanks!

OpenStudy (anonymous):

you're welcome!